OpenStudy (anonymous):

which of the following are equal to the expression below ? Check all that apply sqrt40/2 a. 20 b.sqrt 40/4 c.sqrt 10 d.sqrt 20 ? help plz??

OpenStudy (cwrw238):

sqrt40/2 = 2sqrt10 / 2 = sqrt 10 c

OpenStudy (cwrw238):

need any more explanation?

OpenStudy (cwrw238):

sqrt40/2 = sqrt4 * sqrt10 / 2 = 2 * sqrt10 / 2